If f(x)=3x+1 and g(x)=2x-4, then find (f+g)(0).
Answer by jsmallt9(3758)

The key to this is that (f+g)(x) is short for f(x) + g(x). So in your problem (f+g)(x) = f(x) + g(x) = (3x + 1) + (2x - 4) = 5x - 3

So (f+g)(0) = 5(0) - 3 = 0 - 3 = -3